OpenStudy (anonymous):

Which function below is the inverse of f(x) = The quantity of seven x minus one, over two.?

OpenStudy (08surya):

can u write equation?

OpenStudy (anonymous):

\[f(x)=\frac{ 7x-1 }{ 2 }\]

OpenStudy (anonymous):

i'm about to post the answer choices now.

OpenStudy (08surya):

assume f(x)=y and solve for value of x then change x to y and y to x

OpenStudy (anonymous):

A) \[f ^{-1}(x)=\frac{ 2 }{ 7x+1 }\] B) \[f ^{-1}(x)=\frac{ 2 }{ 1-7x }\] C) \[f ^{-1}(x)=\frac{ 2x+1 }{ 7 }\] D) \[f _{^{-1}}(X)=\frac{ 2x-1 }{ 7 }\]

OpenStudy (anonymous):

that was super confusing

OpenStudy (gracefulchibi):

c

OpenStudy (anonymous):

thank you @gracefulchibi

OpenStudy (08surya):

c

OpenStudy (gracefulchibi):

your welcome :)